OpenStudy (dls):

Approximating 10^(-0.83) =?

OpenStudy (unklerhaukus):

o

OpenStudy (dls):

no :/

OpenStudy (unklerhaukus):

almost

OpenStudy (dls):

\[\LARGE 10^{-6.83}\] I have to find for this...

OpenStudy (dls):

\[\LARGE \frac{10^{-6}}{10^{0.83}}\] Did it this way

OpenStudy (shubhamsrg):

it'll be slightly greater than 0.1

OpenStudy (shubhamsrg):

10^(-0.83) i.e.

OpenStudy (shubhamsrg):

sab Q-tiyapa hai :)

OpenStudy (anonymous):

You could use series expansions, but you would have to calculate (-0.83)^n, so if n is even, there is a complex number. \[10^{x}=\sum_{n=0}^{\infty} \frac{x^{n}\ln^{n}(10)}{n!}\] There is the linear approximation: \[L(x)=f(a)+f'(a)(x-a)\] the derivative of 10^x is ln(x)*10^x, so for negative numbers you would end in complex numbers again. These are the only ones I know. (Sorry for my english)

OpenStudy (anonymous):

complicated question.